A student weighs 0.3800 g of MgCO3. How many moles does this mass represent?
1
Expert's answer
2015-11-05T09:35:46-0500
n = m/M
where n – is a number of moles of a substance(mol), m – mass of a substance(g), M – molar mass of a substance (g/mol).
Then n = 0.38/84.3139 = 0.0045 moles

Answer: 0.0045 moles is the number of moles of MgCO3